# Project Page: https://subjectdrive.github.io/

### v1

@@ 4.2 Implementation Details
To construct the subject bank, we draw from two distinct sources, categorizing them into internal and external subject banks. The internal subject bank is curated by collecting subjects from the training set of the nuScenes dataset. The external subject bank is established by integrating external vehicle datasets from the open-source CompCars [45] dataset. The internal subject bank is used during the training phase, while the external subject bank is mixed with the internal one during the sampling phase to promote the generation of diverse data.
4.3 Main Results
Analysis of Synthetic Data Scale-up. We investigate the impact of scaling up the quantity of generative data on the performance of downstream perception tasks and present the results in Table 1. Two interesting findings emerge. First, increasing the quantity of synthetic data has a positive influence on the performance of the perception model. Although using a small amount of synthetic data is initially less effective than using real nuImages data, increasing the data volume can boost the models to achieve superior performance in both NDS and AMOTA. This demonstrates that scaling up generative data is essential to unlock the potential of generative data production. Second, the integration of external subjects can substantially improve scaling performance. For instance, while tripling the generative data from Panacea yields a mere 0.2 improvement in NDS, incorporating SubjectDrive at the same data scale elevates the NDS by 1.0. Similar trends are also observed in AMOTA. These results demonstrate that utilizing external subjects is an effective way to enhance the generative model’s scaling capability.
Analysis of Performance in 3D Object Detection Task. Next, we present a quantitative analysis that compares SubjectDrive with other data generation methods in the 3D detection task. We utilize SubjectDrive to generate novel synthetic data as additional training source for the StreamPETR model, and then evaluate its performance on the real nuScenes validation set. The results are showcased in Table 2. When trained solely on the synthetic data, SubjectDrive manages to outperform Panacea by 5.5 mAP and 5.0 NDS. Specifically, SubjectDrive achieves an NDS of 41.1, reaching 88% of the performance compared to that trained with real nuScenes data. Furthermore, when combining synthetic data with real data, SubjectDrive surpasses all other methods by a significant margin.
Table 2: Comparison of performance on the 3D object detection task with other generation methods. * indicates the evaluation of WoVoGen is only on the vehicle classes of cars, trucks, and buses.
@@ 4.3 Main Results
Real Real Only 34.5 46.9 59.4 29.1
Generated Panacea[43] 22.5 36.1 72.7 46.9
SubjectDrive 28.0 (+5.5%) 41.1 (+5.0%) 69.5 37.0
Real+Generated DriveDreamer [41] 35.8 39.5 - -
WoVoGen* [27] 36.2 18.1 15.7 123.4
MagicDrive [11] 35.4 39.8 - -
Panacea [43] 37.1 49.2 54.2 27.3
SubjectDrive 38.1 50.2 52.2 26.4
@@
Table 3: Comparison of performance on the 3D object tracking task.
DataType Method AMOTA $\uparrow$ MOTA $\uparrow$ RECALL $\uparrow$ AMOTP $\downarrow$
Real Real Only 30.1 27.1 41.8 1.379
Generated SubjectDrive 23.4 23.5 32.3 1.544
Real+Generated Panacea [43] 33.7 30.6 44.8 1.353
SubjectDrive 37.2 (+3.5%) 33.3 (+2.7%) 47.3 (+2.5%) 1.317
@@
BEVGen[38] $\checkmark$ - 25.54
BEVControl[44] $\checkmark$ - 24.85
MagicDrive[11] $\checkmark$ - 16.20
DriveDreamer[41] $\checkmark$ 452 52.6
WoVoGen [27] $\checkmark$ $\checkmark$ 418 27.6
Panacea [43] $\checkmark$ $\checkmark$ 139 16.96
SubjectDrive $\checkmark$ $\checkmark$ 124 15.98
4.4 Ablation Studies
The alignment between generated samples and the provided BEV conditional labels is essential for evaluating generative models’ controllability. It also acts as a key indicator of the synthetic data’s applicability. In this section, we will conduct ablation studies to justify the design choices in SubjectDrive using this metric.
@@ 4.4 Ablation Studies
[…]
-- -- -- 18.8 32.1 11.4 21.3
$\checkmark$ -- -- 20.0 34.3 15.6 24.5
$\checkmark$ $\checkmark$ -- 22.1 36.1 17.4 27.1
$\checkmark$ $\checkmark$ $\checkmark$ 22.8 (+4.0%) 36.3 (+4.2%) 18.1 (+6.7%) 28.1 (+6.8%)
Real Validation Data 34.5 46.9 30.1 41.8
Table 5 presents our detailed evaluation results for the Subject Visual Adapter (SVA), Subject Prompt Adapter (SPA), and Augmented Temporal Attention (ATA). First, compared to the baseline, introducing the SVA resulted in a significant improvement of 2.2 in NDS and 4.2 in AMOTA, demonstrating its effectiveness in enhancing label alignment. This improvement is achieved by injecting extra subject control through visual features and their corresponding positional embeddings. Second, by integrating the SPA with the SVA, we observed an additional increase of 2.1 in mAP and 1.8 in AMOTA. These results indicate that these two subject control modules are complementary, with each playing a crucial role in the SubjectDrive framework. Third, the final model includes ATA, complementing both SVA and SPA, and yields further improvements of 0.7 in AMOTA and 1.0 in RECALL. These results underscore the efficacy of ATA in enhancing alignment accuracy and temporal consistency across tasks.
4.5 Visualisation results
Figure 3: Generate subject-controlled videos by SubjectDrive.
@@ 4.5 Visualisation results
Figure 4: Multi-view videos generated by SubjectDrive.
@@
Figure 5: Controllable multi-view videos generated by SubjectDrive.
5 Conclusion
In this work, we introduce SubjectDrive, a novel video generation framework designed to enhance the scalability and sampling diversity of generative models. The architecture of SubjectDrive features three key innovations: a subject prompt adapter, a subject visual adapter, and augmented temporal attention. Collectively, these innovations enable a robust subject control capability. Empirically, we find that this subject control feature significantly enhances our model’s ability to produce diverse samples. Through extensive experimentation, SubjectDrive not only demonstrates superior performance but also exhibits strong scaling capabilities. Impressively, our model is the first generative approach to improve the performance of perception models beyond the capabilities of pre-trained models on the nuImages dataset. These results underscore the transformative potential of generative data in advancing autonomous driving technologies, pointing to a promising direction for future developments in the field.

### v2

@@ Introduction
Table 1: Evaluation of data scaling on detection and tracking tasks. Extras denote extra samples beyond nuScenes dataset.
Analysis of Performance in 3D Object Detection Task. Next, we present a quantitative analysis that compares SubjectDrive with other data generation methods in the 3D detection task. We utilize SubjectDrive to generate novel synthetic data as additional training source for the StreamPETR model, and then evaluate its performance on the real nuScenes validation set. The results are showcased in Table 2. When trained solely on the synthetic data, SubjectDrive manages to outperform Panacea by 5.5 mAP and 5.0 NDS. Specifically, SubjectDrive achieves an NDS of 41.1, reaching 88% of the performance compared to that trained with real nuScenes data. When combining synthetic data with real data, SubjectDrive surpasses all other methods by a significant margin.
@@
Real Real Only 34.5 46.9 29.1
Gen Panacea 22.5 36.1 46.9
Ours 28.0 (+5.5%) 41.1 (+5.0%) 37.0
Real+Gen DriveDreamer 35.8 39.5 -
WoVoGen* 36.2 18.1 123.4
MagicDrive 35.4 39.8 -
Panacea 37.1 49.2 27.3
Ours 38.1 50.2 26.4
Table 2: Comparison on the 3D object detection task with other generation methods. * indicates the evaluation of WoVoGen is only on the vehicle classes of cars, trucks, and buses.
@@
DataType Method AMOTA $\uparrow$ AMOTP $\downarrow$ MOTA $\uparrow$
Real Real Only 30.1 1.379 27.1
Gen Ours 23.4 1.544 23.5
Real+Gen Panacea 33.7 1.353 30.6
Ours 37.2 (+3.5%) 1.317 33.3 (+2.7%)
Table 3: Comparison on the 3D object tracking task.
[…]
BEVGen $\checkmark$ - 25.54
BEVControl $\checkmark$ - 24.85
MagicDrive $\checkmark$ - 16.20
DriveDreamer $\checkmark$ 353 26.8
WoVoGen $\checkmark$ $\checkmark$ 418 27.6
Panacea $\checkmark$ $\checkmark$ 139 16.96
Ours $\checkmark$ $\checkmark$ 124 15.98
@@
Ablation Studies
The alignment between generated samples and the provided BEV conditional labels is essential for evaluating generative models’ controllability. It also acts as a key indicator of the synthetic data’s applicability. In this section, we conduct ablation studies to justify the design choices in SubjectDrive using this metric. Table 5 presents our evaluation results for the Subject Visual Adapter (SVA), Subject Prompt Adapter (SPA), and Augmented Temporal Attention (ATA). First, compared to the baseline, introducing the SVA resulted in a significant improvement of 2.2 in NDS and 4.2 in AMOTA, demonstrating its effectiveness in enhancing label alignment. This improvement is achieved by injecting extra subject control through visual features and their corresponding positional embeddings. Second, by integrating the SPA with the SVA, we observed an additional increase of 1.8 in NDS and AMOTA. These results indicate that these two subject control modules are complementary, with each playing a crucial role in the SubjectDrive framework. Third, the final model includes ATA, complementing both SVA and SPA, and yields further improvements of 0.7 in AMOTA and 1.0 in RECALL. These results underscore the efficacy of ATA in enhancing alignment accuracy and temporal consistency across tasks.
@@
-- -- -- 32.1 11.4 21.3
$\checkmark$ -- -- 34.3 15.6 24.5
$\checkmark$ $\checkmark$ -- 36.1 17.4 27.1
$\checkmark$ $\checkmark$ $\checkmark$ 36.3 (+4.2%) 18.1 (+6.7%) 28.1 (+6.8%)
Real Validation 46.9 30.1 41.8
@@
Visualisation results
@@
Figure 7: Multi-view videos generated by Ours.
@@
Figure 8: Subject-controlled videos generated by Ours.
Figure 9: Controllable multi-view videos generated by Ours.
@@
Conclusion
In this work, we present SubjectDrive, a novel video generation framework that enhances the scalability and sampling diversity of generative models. The architecture incorporates three key innovations: a subject prompt adapter, a subject visual adapter, and augmented temporal attention, collectively enabling robust subject control. This feature significantly improves the model’s ability to generate diverse samples. Extensive experiments demonstrate that SubjectDrive not only outperforms existing methods but also scales effectively. Notably, it is the first generative model to enhance perception performance beyond pre-trained capabilities on the nuImages dataset, showcasing the transformative potential of generative data in advancing autonomous driving technologies and pointing to promising future directions in the field.
Acknowledgments
The work was supported by National Science and Technology Major Project of China (2023ZD0121300).